OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x + 1 and whose width is x + 11? Answer x^2+ 12x + 11 x^2+ 11x + 12 x^2+ 10x + 11 x^2+ 11

OpenStudy (hoblos):

area of rectangle = length x width (x + 1)(x + 11) = x^2+ 12x + 11

OpenStudy (anonymous):

Area of a rectangle = lw a = (x+1)(x+11) = x^2+11x+x+11 = x^2+12x+11 area = a: x^2+12x+11